# Synthetic Text Generation for Training Large Language Models via Gradient Matching

### v1

@@ Abstract
Synthetic data has the potential to improve the performance, training efficiency, and privacy of real training examples. Nevertheless, existing approaches for synthetic text generation are mostly heuristics and cannot generate human-readable text without compromising the privacy of real data, or provide performance guarantees for training Large Language Models (LLMs). In this work, we propose the first theoretically rigorous approach for generating synthetic human-readable text that guarantees the convergence and performance of LLMs during fine-tuning on a target task. To do so, we leverage Alternating Direction Method of Multipliers (ADMM) that iteratively optimizes the embeddings of synthetic examples to match the gradient of the target training or validation data, and maps them to a sequence of text tokens with low perplexity. In doing so, the generated synthetic text can guarantee convergence of the model to a close neighborhood of the solution obtained by fine-tuning on real data. Experiments on various classification tasks confirm the effectiveness of our proposed approach.
@@ 1 Introduction
Existing approaches for synthetic text generation mostly rely on advanced LLMs such as GPT-4 to generate synthetic text for the target categories (Ye et al., 2022; Meng et al., 2022; Li et al., 2023b; Gupta et al., 2023; Tao et al., 2024; Wu et al., 2024; Dekoninck et al., 2024; Yu et al., 2024). LLM-generated text either suffers from lack of diversity and faithfulness to real data (Ye et al., 2022; Meng et al., 2022; Li et al., 2023b), or requires meticulous prompt engineering and highly complex pipelines, such as multi-agent frameworks, iterative sampling, and processing mechanisms (Gupta et al., 2023; Dekoninck et al., 2024; Wu et al., 2024). The complexity of the pipelines, efforts for manual prompt engineering, and the cost of querying advanced models limits the applicability of such approaches. A few recent studies explored the use of VAEs and diffusion for controllable text generation (Li et al., 2022; Gong et al., 2022; Zhou et al., 2024). But, training diffusion models is computationally heavy and difficult in practice. Importantly, none of the above approaches provide any guarantee for the performance of the LLM trained on the synthetic text.
@@
The above limitations raise a key question: Can we generate a small subset of synthetic text that can train an LLM with similar dynamics to that of real data? For vision models, Dataset Distillation (DD) addresses the above question by generating a small number of synthetic images that minimize the training loss (Wang et al., 2018; Loo et al., 2022; Nguyen et al., 2020), match the training gradient (Zhao et al., 2020; Zhao & Bilen, 2021) or model’s weight trajectory during training (Cazenavette et al., 2022; Wang et al., 2022). For images, gradient-based methods can easily operate in the pixel-wise continuous space. However, for LLMs, the discrete nature of text and the very large number of LLM’s parameters make DD much more challenging. The few existing approaches generate synthetic embeddings that minimize the training loss (Sucholutsky & Schonlau, 2021; Li & Li, 2021; Sahni & Patel, 2023; Maekawa et al., 2023), or by training a generator model to match the gradient of an LLM trained on the data (Maekawa et al., 2024). However, the synthetic embeddings are not readable and cannot be transferred to train other LLMs, and synthetic data generated by matching dynamics of a model trained on real data may include real training examples and is not privacy preserving.
@@
In this work, we propose the first theoretically-rigorous method to generate readable synthetic text that guarantees similar dynamics to that of fine-tuning on real data. First, we formulate a discrete optimization problem to find text embeddings that have a similar gradient to that of real data, under the constraint that the optimized embeddings should correspond to tokens in the vocabulary. Moreover, to ensure readability, we add another constraint that requires the sequence to have a low perplexity. Then, we solve this discrete optimization problem using Alternating Direction Method of Multipliers (ADMM) that iteratively optimizes the embeddings of synthetic data to match the gradient of the real data, and maps them to a sequence of text tokens with low perplexity. We prove that the synthetic text generated by our method guarantees convergence to a close neighborhood of the solution obtained by fine-tuning the model on real data.
@@
We conduct extensive experiments to evaluate the effectiveness of our approach, namely GrADmm, for generating synthetic data using Phi for multiple classification tasks. First, we consider the case where only a small number of validation examples are available and we apply GrADmm to generate a larger fine-tuning data. We show that with only 5 to 50 examples, GrADmm can successfully generate 100 synthetic data that outperform training on the real examples by up to 32.4%. Next, we apply GrADmm to generate a small synthetic data based on an existing fine-tuning data. We show that the synthetic data generated by GrADmm outperforms zero-shot and few-shot generation by LLMs as well as real examples selected by coreset selection methods by up to 10.4%, while ensuring the privacy of the training data. We also confirm the transferability of GrADmm’s generated text via Phi for fine-tuning other LLMs, including Llama-3.2-1B and OPT-1.3B.
@@ 2.1 Dataset Distillation (DD)
DD aims to generate a small synthetic subset of examples that can achieve a similar generalization performance to that of training on the full real dataset. DD is one of the most effective ways to preserve the privacy of training data.
DD for Images. DD is originally proposed for images. For image datasets, Wang et al. (2018) originally proposed a meta-learning approach which synthesizes data by iteratively training a model to convergence on the synthetic examples, and optimizing the synthetic data such that the trained model generalizes well on the real training data. Subsequent studies tried to make this process more efficient by using kernel methods to approximate training the model on synthetic data in a closed form (Loo et al., 2022; Nguyen et al., 2020). More recent works generate synthetic data by matching the gradient (Zhao et al., 2020; Zhao & Bilen, 2021; Kim et al., 2022) or wright trajectory (Cazenavette et al., 2022; Wang et al., 2022) of the model trained on real data, or by matching the data distribution (Zhao & Bilen, 2023).
DD for Text. There has been recent efforts in applying DD to text. For text datasets, existing methods (Sucholutsky & Schonlau, 2021; Li & Li, 2021; Sahni & Patel, 2023) apply the original meta-learning based method of (Wang et al., 2018), or minimize the KL-divergence between the self-attention probabilities of the model and the distilled attention labels across all layers and heads, for the first token Maekawa et al. (2023). As generating text in the discrete space is difficult, the synthetic data is generated as continuous input word embeddings instead of discrete text. Such embeddings cannot be used for training other models that have different word embedding weights, and are unreadable to humans, making them difficult to interpret and analyze. Sucholutsky & Schonlau (2021); Sahni & Patel (2023) transformed their distilled synthetic samples to text by finding words that has the nearest neighbor embeddings. However, this results in unrelated words that are not meaningful.
To generate readable text, Maekawa et al. (2024) first trains a proxy language model from scratch to generate synthetic training data for different classes. Then, it fine-tunes a generator model to generate synthetic data by minimizing the gradient matching loss between generated and training data. Training the proxy model is a bottleneck in scaling the method. Besides, as the distilled synthetic data may include real samples from the original dataset, this method does not have the privacy-preserving advantage of prior DD methods.
@@ 4.2 Alternating Between Text and Embedding Spaces
Constrained Gradient Matching in the Embedding Space. To apply ADMM to our discretely constrained non-convex problem 4, we convert it to a non-convex optimization with convex linear constraints. To do so, we introduce an auxiliary variable $\boldsymbol{Z}$ and rewrite our objective with an extra equality constraint so that the embeddings are constrained to be from the vocabulary, but not subject to that restriction:
@@
The indicator function ${\mathcal{I}}_{\mathcal{E}}(\boldsymbol{Z})$ is defined as ${\mathcal{I}}_{\mathcal{E}}(\boldsymbol{Z})=0$ if ${z}_{j}\in{\mathcal{E}}~{}\forall j$ (i.e., if the embedding of each synthetic example can be mapped to a sequence of words in the vocabulary), and ${\mathcal{I}}_{\mathcal{E}}(\boldsymbol{Z})=+\infty$ otherwise.
The augmented Lagrange of Eq. 5 for parameter $\rho>0$, can be formulated as:
@@
Projecting the Embeddings into the Vocabulary Space. The projection step in Eq. 9 can be written as (Huang et al., 2021):
@@
1:  Input: Constant $\rho>0$, ADMM steps $T$, proj param $k$
@@
4:  Initialize $\boldsymbol{X}^{0}=\operatorname*{arg\,min}_{\boldsymbol{X}}f(\boldsymbol{X})$
@@
8:     Update $\boldsymbol{X}$: $\boldsymbol{X}^{t+1}=\operatorname*{arg\,min}_{\boldsymbol{X}}{\mathcal{L}}(\boldsymbol{X},\boldsymbol{Z}^{t},\boldsymbol{\Lambda}^{t})$
@@ 4.3 Dealing with High-dimension Gradients
Figure 1: Data-scarce regime. Generating 100 synthetic samples with GrADmm, based on 5, 10, 20, 50 examples from a target task. Synthetic data generated based on only 5 real examples outperforms the real data by 18.6%, 28.5%, and 32.4% on the three datasets.
@@ 4.5 Convergence Analysis
Thus, the fine-tuned models have a similar performance.
@@
Dataset # data acc ppl acc ppl acc ppl acc ppl acc ppl acc ppl 1K
5 87.0 5.8 73.2 2.5 72.4 3.0 68.0 6.6 76.6 5.5 53.6 7.7
SST-2 10 88.1 5.1 80.7 2.3 71.8 2.6 79.9 6.7 82.2 5.5 62.7 7.3
(classification) 20 88.3 5.3 84.1 2.2 76.8 2.7 87.3 6.6 87.9 5.7 86.3 7.0 91.2
50 89.4 5.7 84.1 2.3 75.0 2.6 87.5 6.6 88.6 5.5 87.0 6.8
pretrained: 69.6 100 90.0 5.2 88.0 2.3 76.4 2.5 89.0 6.7 89.6 5.8 89.7 6.7
Tweet 5 84.2 3.5 78.1 2.7 54.6 3.0 56.5 5.8 56.5 5.1 56.5 5.9
emotions 10 84.3 2.8 78.8 2.3 47.8 3.2 58.8 5.6 70.1 5.6 62.4 5.8
(classification) 20 85.7 3.4 82.9 2.2 70.7 3.3 70.3 5.8 79.3 5.5 72.1 5.8 96.3
50 86.1 4.3 83.9 2.2 80.0 3.1 79.5 5.6 85.0 5.1 77.8 5.6
pretrained: 43.7 100 86.4 3.8 84.8 2.3 84.5 3.4 86.1 5.5 86.0 5.1 83.4 5.5
Rotten 5 82.6 4.5 72.7 2.5 59.5 3.8 73.4 4.9 61.5 6.4 73.7 5.4
tomatoes 10 83.1 5.5 75.9 2.3 66.7 3.0 74.3 5.2 63.8 5.5 74.9 5.8
(classification) 20 84.8 7.0 78.5 2.2 78.5 3.1 80.6 5.7 68.4 5.2 79.0 5.7 88.4
50 85.2 4.6 77.7 2.3 79.9 2.9 81.6 5.6 80.0 5.1 81.8 5.6
pretrained: 65.8 100 85.2 4.5 82.2 2.3 82.3 2.9 84.0 5.6 83.5 5.1 84.3 5.6

### v2

@@ Abstract
Synthetic data has the potential to improve the performance, training efficiency, and privacy of real training examples. Nevertheless, existing approaches for synthetic text generation are mostly heuristics and cannot generate human-readable text without compromising the privacy of real data, or provide performance guarantees for training Large Language Models (LLMs). In this work, we propose the first theoretically rigorous approach for generating synthetic human-readable text that provides convergence, performance, and privacy guarantees for fine-tuning LLMs on a target task. To do so, we leverage Alternating Direction Method of Multipliers (ADMM) that iteratively optimizes the embeddings of synthetic examples to match the noisy gradient of the target training or validation data, and maps them to a sequence of text tokens with low perplexity. In doing so, the generated synthetic text guarantees convergence of the model to a close neighborhood of the solution obtained by fine-tuning on real data and preserves their privacy. Experiments on various classification tasks confirm the effectiveness of our proposed approach. Our code is available at https://github.com/BigML-CS-UCLA/GRADMM.
@@ 1 Introduction
Existing approaches for synthetic text generation mostly rely on advanced LLMs such as GPT-4 to generate synthetic text for the target categories (Ye et al., 2022; Meng et al., 2022; Li et al., 2023b; Gupta et al., 2023; Tao et al., 2024; Wu et al., 2024; Dekoninck et al., 2024; Yu et al., 2024). LLM-generated text either suffers from lack of diversity and faithfulness to real data (Ye et al., 2022; Meng et al., 2022; Li et al., 2023b), or requires meticulous prompt engineering and highly complex pipelines, such as multi-agent frameworks, iterative sampling, and processing mechanisms (Gupta et al., 2023; Dekoninck et al., 2024; Wu et al., 2024). The complexity of the pipelines, efforts for manual prompt engineering, and the cost of querying advanced models limits the applicability of such approaches. A few recent studies explored the use of VAEs and diffusion for controllable text generation (Li et al., 2022; Gong et al., 2022; Zhou et al., 2024). But, training diffusion models is computationally heavy and difficult in practice. Importantly, none of the above approaches guarantees performance of LLMs trained on the synthetic text or preserve privacy of real data.
@@
The above limitations raise a key question: Can we generate a small subset of synthetic text that can train an LLM with similar dynamics to that of real data? For vision models, Dataset Distillation (DD) addresses the above question by generating a small number of synthetic images that minimize the training loss (Wang et al., 2018; Loo et al., 2022; Nguyen et al., 2020), match the training gradient (Zhao et al., 2020; Zhao & Bilen, 2021) or model’s weight trajectory during training (Cazenavette et al., 2022; Wang et al., 2022). For images, gradient-based methods can easily operate in the pixel-wise continuous space. However, for LLMs, the discrete nature of text and the very large number of LLM’s parameters make DD much more challenging. The few existing approaches generate synthetic embeddings that minimize the training loss (Sucholutsky & Schonlau, 2021; Li & Li, 2021; Sahni & Patel, 2023; Maekawa et al., 2023), or by training a generator model to match the gradient of an LLM trained on target data (Maekawa et al., 2024). However, the synthetic embeddings are not readable and cannot be transferred to train other LLMs, and synthetic data generated by matching dynamics of a LLM trained on target data may include real training examples and is not privacy preserving.
@@
In this work, we propose the first theoretically-rigorous method to generate readable synthetic text that guarantees similar dynamics to that of fine-tuning on real data. First, we formulate a discrete optimization problem to find text embeddings that have a similar gradient to that of real data, under the constraint that the optimized embeddings should correspond to tokens in the vocabulary. Moreover, to ensure readability, we add another constraint that requires the sequence to have a low perplexity. Then, we solve this discrete optimization problem using Alternating Direction Method of Multipliers (ADMM) that iteratively optimizes the embeddings of synthetic data to match the average gradient of the real data, and maps them to a sequence of text tokens with low perplexity. To guarantee Differential Privacy (DP), we clip real data gradients and add controlled noise to their average before matching it. We prove that the synthetic text generated by our method guarantees convergence to a close neighborhood of the solution obtained by fine-tuning the model on real data.
@@
We conduct extensive experiments to evaluate the effectiveness of our approach, namely GrADmm, for generating synthetic data using Phi model for multiple classification tasks. First, we consider the case where only a small number of validation examples are available and we apply GrADmm to generate a larger fine-tuning data. We show that with only 5 to 50 examples, GrADmm can successfully generate 100 synthetic data that outperform training on the real examples by up to 31.5%. Next, we apply GrADmm to generate a small synthetic data based on an existing larger fine-tuning data. We show that the synthetic data generated by GrADmm outperforms zero-shot and few-shot generation by LLMs as well as real examples selected by coreset selection methods by up to 13.1%, while ensuring the privacy of the training data. We also confirm the transferability of GrADmm’s generated text via Phi for fine-tuning other LLMs, including Llama-3.2-1B and OPT-1.3B.
@@ 2.1 Dataset Distillation (DD)
DD aims to generate a small synthetic subset of examples that can achieve a similar generalization performance to that of training on the full real dataset.
DD for Images. DD is originally proposed for images. Wang et al. (2018) initially proposed a meta-learning approach which synthesizes data by iteratively training a model to convergence on the synthetic examples, and optimizing the synthetic data such that the trained model generalizes well on the real training data. Subsequent studies tried to make this process more efficient by using kernel methods to approximate training the model on synthetic data in a closed form (Loo et al., 2022; Nguyen et al., 2020). More recent works generate synthetic data by matching the gradient (Zhao et al., 2020; Zhao & Bilen, 2021; Kim et al., 2022) or wright trajectory (Cazenavette et al., 2022; Wang et al., 2022) of the model trained on real data, or by matching the data distribution (Zhao & Bilen, 2023).
DD for Text. There have been recent efforts in applying DD to text. For text datasets, existing methods (Sucholutsky & Schonlau, 2021; Li & Li, 2021; Sahni & Patel, 2023) apply the original meta-learning based method of (Wang et al., 2018), or minimize the KL-divergence between the self-attention probabilities of the model and the distilled attention labels across all layers and heads, for the first token Maekawa et al. (2023). As generating text in the discrete space is difficult, the synthetic data is generated as continuous input word embeddings instead of discrete text. Such embeddings cannot be used for training other models that have different word embedding weights, and are unreadable to humans, making them difficult to interpret and analyze. Sucholutsky & Schonlau (2021); Sahni & Patel (2023) transformed their distilled synthetic samples to text by finding words with the nearest neighbor embeddings. However, this results in unrelated words that are not meaningful.
To generate readable text, Maekawa et al. (2024) first trains a proxy language model from scratch to generate synthetic training data for different classes. Then, it fine-tunes a generator model to generate synthetic data by minimizing the gradient matching loss between generated and training data. Training the proxy model is a bottleneck in scaling the method. Besides, as the distilled synthetic data may include real samples from the original dataset, this method cannot ensure privacy.
@@ 4.2 Alternating Between Text and Embedding Spaces
Constrained Gradient Matching in the Embedding Space.
To apply ADMM to our discretely constrained non-convex problem 4, we convert it to a non-convex optimization with convex linear constraints. To do so, we introduce an auxiliary variable $\boldsymbol{Z}$ and rewrite our objective with an extra equality constraint so that the embeddings are constrained to be from the vocabulary, but not subject to that restriction:
@@
The indicator function ${\mathcal{I}}_{\mathcal{E}}(\boldsymbol{Z})$ is defined as ${\mathcal{I}}_{\mathcal{E}}(\boldsymbol{Z})=0$ if ${z}_{j}\in{\mathcal{E}}~{}\forall j$ (i.e., if the embedding of each synthetic example can be mapped to a sequence of words in the vocabulary), and ${\mathcal{I}}_{\mathcal{E}}(\boldsymbol{Z})=+\infty$ otherwise. The augmented Lagrange of Eq. 5 for parameter $\rho>0$, can be formulated as:
@@
1:  Input: Constant $\rho>0$, ADMM steps $T$, proj param $k$, DP param $\varepsilon,\delta$
@@
4:  Initialize $\boldsymbol{X}^{0}=\operatorname*{arg\,min}_{\boldsymbol{X}}f(\boldsymbol{X},{\varepsilon,\delta})$
@@
8:     Update $\boldsymbol{X}$: $\boldsymbol{X}^{t+1}=\operatorname*{arg\,min}_{\boldsymbol{X}}{\mathcal{L}}(\boldsymbol{X},\boldsymbol{Z}^{t},\boldsymbol{\Lambda}^{t},{\varepsilon,\delta})$
@@
Projecting the Embeddings into the Vocabulary Space. Equation 9 can be written as (Huang et al., 2021):
@@ 4.3 Dealing with High-dimension Gradients
Figure 1: Data-scarce regime. Generating 100 synthetic samples with GrADmm, based on 5, 10, 20, 50 examples from a target task. Synthetic data generated based on only 5 real examples outperforms the real data by 15.7%, 29.7%, and 31.5% on the three datasets.
@@ 4.4 Filtering the Generated Examples
Remark. Due to top-$k$ decoding, the filtered synthetic examples do not match the target gradient very accurately. Thus, we generate each synthetic example to match the target gradient independently, not conditioned on each other. We will confirm in our experiments in Sec. 5.3 that the synthetic data generated independently by GrADmm matches the real data gradient closely during fine-tuning.
4.5 Making GrADmm Differentially Private
Differential privacy (DP) (Dwork et al., 2006) is a rigorous mathematical framework that ensures no single data point can be identified or inferred from the output of a statistical or machine learning model. To make GrADmm differentially private, we inject controlled noise $\alpha$ into the clipped gradient of the real data in Equation 4. Specifically, GrADmm first computes per-sample gradients of the real data and clips their $\ell_{2}$-norm to a threshold of $C$. These clipped gradients are then averaged, and Gaussian noise, drawn from $\mathcal{N}(0,\sigma^{2})$, is added to this average. The added noise scale $\sigma$ is defined as follows:
[…]
where $c=\sqrt{\log\left(\frac{2}{\sqrt{16\delta+1}-1}\right)}$ and the clipping threshold $C=1$ for all experiments. Based on the composition theorem (Dwork et al., 2010), GrADmm achieves $(\varepsilon,\delta)$-DP.
We denote the new optimization problem and its augmented Lagrangian objective as $f(X,\varepsilon,\delta)$ and ${\mathcal{L}}(\boldsymbol{X},\boldsymbol{Z},\boldsymbol{\Lambda},\varepsilon,\delta)$, respectively. As $\varepsilon\to\infty$, the privacy constraint is relaxed and $f(X,\varepsilon,\delta)\to f(X)$, yielding the original optimization problem. The new problem retains the same structure and can be solved using the ADMM procedure described before.
@@ 4.6 Convergence Analysis
Thus, the fine-tuned models will have a similar performance.
@@
Dataset # data $\epsilon=\infty$ $\epsilon=0.05$ ppl acc ppl acc ppl acc ppl acc ppl acc ppl 1K
5 $\textbf{86.5}_{\pm 0.5}$ $84.2_{\pm 0.1}$ 5.8 $71.6_{\pm 4.4}$ 2.5 $71.8_{\pm 0.7}$ 3.0 $61.2_{\pm 5.7}$ 6.6 $75.3_{\pm 0.9}$ 5.5 $52.9_{\pm 2.4}$ 7.7
SST-2 10 $\textbf{87.4}_{\pm 0.9}$ $86.2_{\pm 1.8}$ 5.1 $79.1_{\pm 6.5}$ 2.3 $72.0_{\pm 0.3}$ 2.6 $78.8_{\pm 0.9}$ 6.7 $82.0_{\pm 1.0}$ 5.5 $66.1_{\pm 11.5}$ 7.3
20 $\textbf{87.9}_{\pm 0.3}$ ${87.2}_{\pm 0.8}$ 5.3 $82.2_{\pm 2.2}$ 2.2 $77.5_{\pm 0.5}$ 2.7 $83.3_{\pm 2.8}$ 6.6 $86.2_{\pm 0.9}$ 5.7 $86.2_{\pm 1.1}$ 7.0 91.3
Base acc: 50 $\textbf{89.7}_{\pm 0.2}$ $88.0_{\pm 0.1}$ 5.7 $83.0_{\pm 1.1}$ 2.3 $80.6_{\pm 3.9}$ 2.6 $86.0_{\pm 0.4}$ 6.6 $88.1_{\pm 1.0}$ 5.5 $87.8_{\pm 0.9}$ 6.8 ±0.4
69.6% 100 $\textbf{89.7}_{\pm 0.1}$ $88.6_{\pm 0.5}$ 5.2 $87.5_{\pm 0.6}$ 2.3 $77.4_{\pm 1.7}$ 2.5 $88.9_{\pm 0.1}$ 6.7 $89.3_{\pm 0.5}$ 5.8 $88.8_{\pm 0.8}$ 6.7
Tweet 5 $\textbf{83.8}_{\pm 0.3}$ $80.4_{\pm 1.7}$ 3.5 $70.7_{\pm 10.1}$ 2.7 $52.5_{\pm 2.8}$ 3.0 $56.3_{\pm 0.3}$ 5.8 $55.4_{\pm 1.6}$ 5.1 $56.2_{\pm 0.8}$ 5.9
emotions 10 $\textbf{84.1}_{\pm 0.2}$ $81.4_{\pm 3.3}$ 2.8 $70.3_{\pm 11.9}$ 2.3 $47.5_{\pm 0.3}$ 3.2 $58.4_{\pm 0.4}$ 5.6 $61.0_{\pm 6.5}$ 5.6 $62.1_{\pm 1.5}$ 5.8
20 $\textbf{85.2}_{\pm 0.6}$ $83.5_{\pm 1.4}$ 3.4 $81.7_{\pm 1.9}$ 2.2 $68.2_{\pm 3.1}$ 3.3 $65.8_{\pm 3.4}$ 5.8 $73.5_{\pm 4.4}$ 5.5 $70.6_{\pm 4.9}$ 5.8 $96.1$
Base acc: 50 $\textbf{85.8}_{\pm 0.3}$ $83.2_{\pm 2.0}$ 4.3 $82.7_{\pm 1.9}$ 2.2 $79.0_{\pm 2.7}$ 3.1 $76.7_{\pm 2.9}$ 5.6 $83.9_{\pm 1.2}$ 5.1 $77.4_{\pm 3.1}$ 5.6 ±0.2
43.7% 100 $\textbf{86.5}_{\pm 0.1}$ $83.9_{\pm 2.0}$ 3.8 $84.2_{\pm 0.6}$ 2.3 $83.5_{\pm 1.4}$ 3.4 $85.7_{\pm 0.4}$ 5.5 $84.6_{\pm 1.5}$ 5.1 $80.8_{\pm 5.0}$ 5.5
Rotten 5 ${82.2}_{\pm 0.3}$ $\textbf{82.4}_{\pm 0.6}$ 4.5 $72.6_{\pm 2.8}$ 2.5 $55.7_{\pm 2.8}$ 3.8 $69.8_{\pm 3.0}$ 4.9 $60.0_{\pm 1.2}$ 6.4 $70.4_{\pm 4.2}$ 5.4
tomatoes 10 $\textbf{82.9}_{\pm 0.2}$ $81.8_{\pm 0.9}$ 5.5 $75.3_{\pm 2.8}$ 2.3 $63.7_{\pm 2.9}$ 3.0 $71.5_{\pm 2.9}$ 5.2 $61.1_{\pm 3.8}$ 5.5 $74.5_{\pm 4.1}$ 5.8
20 $\textbf{84.4}_{\pm 0.5}$ $83.2_{\pm 0.6}$ 7.0 $78.0_{\pm 0.5}$ 2.2 $75.7_{\pm 2.4}$ 3.1 $79.1_{\pm 1.2}$ 5.7 $67.5_{\pm 0.7}$ 5.2 $80.6_{\pm 0.9}$ 5.7 $88.1$
Base acc: 50 $\textbf{84.9}_{\pm 0.2}$ $83.1_{\pm 0.6}$ 4.6 $77.5_{\pm 0.2}$ 2.3 $78.7_{\pm 0.9}$ 2.9 $81.2_{\pm 0.7}$ 5.6 $78.7_{\pm 1.5}$ 5.1 $81.1_{\pm 1.8}$ 5.6 ±0.3
65.8% 100 $\textbf{85.0}_{\pm 0.3}$ $83.2_{\pm 0.7}$ 4.5 $81.3_{\pm 1.0}$ 2.3 $82.3_{\pm 0.3}$ 2.9 $82.8_{\pm 1.2}$ 5.6 $82.1_{\pm 1.2}$ 5.1 $83.7_{\pm 1.1}$ 5.6